how to do the percent yield in organic chemistry lab and how do you find the theoretical yield?

Also how to do the percent recovery?
Answer #1

theoretical yield - theoretical yield is amount predicted by stochiometric calculation based on no. of mole of all reactant present.

for example take simple reaction

CH3 - CH3 + HNO3   \rightarrow  CH3 - CH2 - NO2 + H2O

in this reaction 1 mole of ethane react with 1 mole of nitric acid to produce 1 mole of nitroethane

molar mass of ethane = 30 gm/mol and molar mass of nitroethane = 75 gm/mol

if you reacted 1mole(30 gm) of ethane with excess amount of nitric acid you will get 1 mole (75 gm ) of nitroethane

this 75 gm of nitroethane is theoritical yield.

percent yield -

percent yield = Actual yield \times 100 / theoritical yield

if we carry out reaction we will not get 100 percent yield that mean theorical yield. then those yield we get calculated by as percent yield

for example in the above reaction of nitration of ethane 30 gm ethane give 75 gm of nitroethane it is theoritical yield

but practcally if we get only 65 gm of product then percent yield = 65 \times 100 / 75 = 86.67 %

percent recovery - percent recovery term used when there is no reaction such as purification of sample.